OpenStudy (anonymous):

Simply and show work (-2+4i)+(7-3i)

OpenStudy (jlg030597):



OpenStudy (jlg030597):

@kc_kennylau Is this right

OpenStudy (jlg030597):

just Checking

OpenStudy (kc_kennylau):

this is not an equation, don't add 3i to both parentheses

OpenStudy (jlg030597):

but you simplify it I thought thats how you would simplify it

OpenStudy (kc_kennylau):

\[(-2+4i)+(7-3i)\\=-2+4i+7-3i\\=-2+7+4i-3i\\=-5+i\]

OpenStudy (jlg030597):

oh ok that makes since now so if there werent ( ) Id be right

OpenStudy (kc_kennylau):

If it were \(-2+4i\color{orange}=7-3i\) you would be right

OpenStudy (kc_kennylau):

actually no since i is not an unknown

OpenStudy (jlg030597):

oh ok now I get it. I need to do a little more studying on math

OpenStudy (kc_kennylau):

\[\Large i=\sqrt{-1}\]

OpenStudy (kc_kennylau):

@jlg030597 good luck with your studying :D

OpenStudy (jlg030597):

Thanks lol Ill see yall later bye

OpenStudy (kc_kennylau):

bye :)